OpenStudy (anonymous):

FAN AND MEDAL: Which is a factored form of 216p^3+125q^3?

OpenStudy (solomonzelman):

At first, you can factor out of q^2.

OpenStudy (anonymous):

a. (6p+5q)(36p^2+30pq+25q) b. (6p+5q)(36p^2-30pq-25q) c. (6p+5q)(36p^2+30pq-25q) d. (6p+5q)(36p^2-30pq-25q)

OpenStudy (solomonzelman):

ohh it is p^3 and q^3, my bad.

OpenStudy (anonymous):

Cant you cube root the whole thing

OpenStudy (solomonzelman):

\(\large\color{blue}{ 5^3=125 }\) \(\large\color{blue}{ 6^3=216 }\)

OpenStudy (solomonzelman):

\(\large\color{blue}{ 125q^3=5^3q^3=(5q)^3 }\) \(\large\color{blue}{ 216p^3=6^3p^3=(6p)^3 }\)

OpenStudy (solomonzelman):

Then apply. \(\large\color{blue}{ a^3+b^3=(a+b)(a^2+ab+b^2) }\)

OpenStudy (solomonzelman):

I would check my rule though, It might be wrong.

OpenStudy (anonymous):

so then it would be a?

OpenStudy (solomonzelman):

I messed up my rule. it should say: \(\large\color{blue}{ a^3+b^3=(a+b)(a^2-ab+b^2)}\)

OpenStudy (solomonzelman):

So it's not A,(as you can see)

OpenStudy (anonymous):

ok so then d

OpenStudy (solomonzelman):

it is not -25q at the end, it is supposed to be +25q. No?

OpenStudy (anonymous):

I think i put the options wrong. my paper has +25q at the end

OpenStudy (solomonzelman):

so with +25q at the end , (then) D would be correct.

OpenStudy (anonymous):

Thank you :)

OpenStudy (solomonzelman):

yw